OpenStudy (anonymous):

The measure of a complement of an angle is three more than twice the measure of the angle. Find the measure of the angle and its complement. How can a complement of an angle be more than the angle? or which angle are they talking about? What am I supposed to do?

OpenStudy (lgbasallote):

let a and b be the complements. since they are complements a + b = 90. okay with that?

OpenStudy (anonymous):

Sum of An Angle and its Compliment is always 90 degrees.. So if angle you are supposing as x and its compliment will be : 90 - x Now make the equation according to the question using this concept..

OpenStudy (anonymous):

Compliment Angle = 3 More Than 2*(Angle) Substitute here the values that we are supposing and solve for x..

OpenStudy (anonymous):

wait the complement angle is the same as the other angle that adds up to 90?

OpenStudy (anonymous):

okay I get the rest, but I just need to know what "angle" the problem is talking about. "find the measures of the angle and it's complement" of the angle, is refering to which one, the 90 angle? or the one that adds up to 90?

OpenStudy (lgbasallote):

i think you misunderstood the problem..

OpenStudy (lgbasallote):

or maybe the concept of complementary angles.. complementary angles are two angles that add up to 90. neither one is 90 degrees

OpenStudy (lgbasallote):

for example 45 and 45 are complementary, 30 and 60 are complementary, 20 and 70 are complementary, etc

OpenStudy (anonymous):

yes I know that. "the measure of a complement of an angle is three more than twice the measure of the angle. "the angle", is referring to what? how come they didn't say "the other angle?" because I assume "the angle" is the angle that is being added up to 90.

OpenStudy (lgbasallote):

the angle could be any of the two. a + b = 90..it could be a, it could be b.

OpenStudy (anonymous):

technically... "the angle" is also a complementary XD... right? so I'm right right..

OpenStudy (lgbasallote):

"the measure of a complement of an angle" <--let's say you take b as the complement of an angle..therefore a will be the angle according to some grammar thingies.

OpenStudy (lgbasallote):

does that make sense lol

OpenStudy (anonymous):

so, b(complement)+a(the angle) = 90. so, 2b+3 +a = 90?

OpenStudy (lgbasallote):

2a not 2b...a is "the angle" remember

OpenStudy (lgbasallote):

and also it's impossible for b = 2b + 3

OpenStudy (anonymous):

either way it could work cause they're being added...

OpenStudy (lgbasallote):

well maybe it is...but that doesnt make much sense

OpenStudy (lgbasallote):

not really...if you look at 2b + 3 + a = 90 you cannot combine anything...it's dead end

OpenStudy (anonymous):

OHWAIT.. 2a+3+a=90

OpenStudy (anonymous):

2a+3 being the complement

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

LOL thanks so being soo patient with me XD. my brain is kinda slow.. so

OpenStudy (lgbasallote):

haha no worries \[\Huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]